OpenStudy (anonymous):

h(g(x)) if h(x)=x^2-4x+7 and g(x)=x+1

OpenStudy (anonymous):

h(g(x)) = (?)^2 -4(?) + 7 fill in the ? with x+1

OpenStudy (anonymous):

so you'll get: h(g(x)) = (x+1)^2 - 4(x+1) + 7

OpenStudy (anonymous):

now simplify... h(g(x)) = (x^2 + 2x + 1) - 4x - 4 + 7 = x^2 - 2x + 8